OpenStudy (anonymous):

Multiply. (–4) ∙ (–8) A. –32 B. –12 C. 12 D. 32

OpenStudy (defiance):

4 times 8.

OpenStudy (alexandervonhumboldt2):

4*8=32 -*-=+ thus -4*-8=4*8=???

OpenStudy (anonymous):

sorry this isnt in the math group.

OpenStudy (anonymous):

32

OpenStudy (defiance):

a negative times a negative is a positive

OpenStudy (defiance):

- 4 times -8 = 32

OpenStudy (richgall____):

32

OpenStudy (anonymous):

so its 32. not negative 32

OpenStudy (richgall____):

yea

OpenStudy (defiance):

@richgall____ @CandyCove Don't give direct answers pls

OpenStudy (defiance):

Yes. negitive times negitive is a positive

OpenStudy (richgall____):

a negative times a negative equals a positive

OpenStudy (richgall____):

-_- dude

OpenStudy (anonymous):

ok thanks you guys. @richgall____ @Defiance @CandyCove

OpenStudy (defiance):

Np.

OpenStudy (richgall____):

welcome.